The owner of a moving company usually has her most experienced manager predict the total number of labor hours required to complete a move. While useful, the owner is interested in a more accurate method of predicting the number of labor hours required. As a preliminary effort, data was collected on the number of work hours required to complete a move, number of cubic feet moved, the number of large pieces, and if an elevator was present.0 eeeeeeee o e e e e e e e o o e e e e e e e oeo EYYNNYYYYYYYYYYYYNYYYYYYYNNYYYYYYYNYN 7 3222132161443232222543323333434325 0

How much correlation is there between the variables?

How much of the variability in strength is explained by the predictors?

Estimate the number of hours required when 325 cubic feet are moved with 3 large pieces of furniture and no elevator is present.

Estimate the number of hours required when 425 cubic feet are moved with 7 large pieces of furniture and an elevator is present.

Estimate the number of hours required when 375 cubic feet are moved with 2 large pieces of furniture and no elevator is present.
Answer #1

excel o/p for regression using data analysis tool g/n below -

SUMMARY OUTPUT
Regression Statistics
Multiple R 0.980588
R Square 0.961552
Adjusted R Square 0.957948
Standard Error 3.05571
Observations 36
ANOVA
df SS MS F Significance F
Regression 3 7472.642 2490.881 266.7649 1.03E-22
Residual 32 298.7956 9.337362
Total 35 7771.438
Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Lower 95.0% Upper 95.0%
Intercept 2.990449 1.907844 1.56745 0.126846 -0.8957 6.876599 -0.8957 6.876599
feet 0.025598 0.003764 6.801469 1.09E-07 0.017932 0.033264 0.017932 0.033264
large 5.042393 0.721561 6.988175 6.43E-08 3.572622 6.512164 3.572622 6.512164
elevator -6.76828 1.382064 -4.89723 2.68E-05 -9.58345 -3.95311 -9.58345 -3.95311

correlation is 0.9806

R Square   0.9616

so, 96.16% of variability in strength is explained by the predictors.

Estimate the number of hours required when 325 cubic feet are moved with 3 large pieces of furniture and no elevator is present

estimated hours = 2.99*feet + 5.042 * large - 6.768*elevator

so, estimated hours = 2.99*325 + 5.042*3 - 6.768*0 = 26.44

----------------------------------------------------------------------

Estimate the number of hours required when 425 cubic feet are moved with 7 large pieces of furniture and an elevator is present.

estimated hours = 2.99*feet + 5.042 * large - 6.768*elevator

estimated hours = 2.99*425 + 5.042 * 7 - 6.768*1 = 42.40

--------------------------------------------------------------------------------------

Estimate the number of hours required when 425 cubic feet are moved with 7 large pieces of furniture and an elevator is present.

estimated hours = 2.99*feet + 5.042 * large - 6.768*elevator

estimated hours = 2.99*375 + 5.042 * 2 - 6.768*0 = 22.67